Below are 36 sorted ages of an acting award winner. Find Upper P 60P60 using the method presented in the textbook. 1818 2121 2121 2222 2323 2525 2727 3030 3131 3131 3434 3636 3838 3939 4444 4646 4646 5050 5252 5252 5454 5757 5757 6363 6464 6565 6565 6565 6565 6666 6868 6868 6868 7272 7777 7878
Answer #1

Solution:- P60 = 57.0

We know that pth percentile =(p*(n+1)th) /100 value

Therefore P60 =60*(36+1)/100 = 22.2 position

Now we need to find the value of 22.2nd position.

22.2nd position = Value of 22nd position + Value of 0.2nd position

22.2nd position = 57 + 0.2* (57 - 57) = 57+ 0

P60 = 57.0